OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. y − 6 = −3x − 3 y − 6 = −3(x + 1) y = −3x + 3 3x + y = 3

OpenStudy (anonymous):

I think its the 2nd one

OpenStudy (anonymous):

yessir it it, you dont have a graphing calculator?

OpenStudy (anonymous):

oh it really is the second one and no i dont have a graphing caculator.

OpenStudy (anonymous):

wait?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

point slope form is y-y1=m(x-x1), given y1= 6 and x1=-1 an your slope m=-3 plug in values

OpenStudy (anonymous):